1/2 + (2/3 / 3/4) - (4/5*5/6)=?
Answer by rfer(16322)

1/2+(2/3/3/4)-(4/5*5/6)=
1/2+(2/3*4/3)-20/30=
1/2+8/9-20/30=
45/90+80/90-60/90=65/90=13/18